### v1

# All-in-One Slider for Attribute Manipulation in Diffusion Models

Abstract
Text-to-image (T2I) diffusion models have made significant strides in generating high-quality images. However, progressively manipulating certain attributes of generated images to meet the desired user expectations remains challenging, particularly for content with rich details, such as human faces. Some studies have attempted to address this by training slider modules. However, they follow a One-for-One manner, where an independent slider is trained for each attribute, requiring additional training whenever a new attribute is introduced. This not only results in parameter redundancy accumulated by sliders but also restricts the flexibility of practical applications and the scalability of attribute manipulation. To address this issue, we introduce the All-in-One Slider, a lightweight module that decomposes the text embedding space into sparse, semantically meaningful attribute directions. Once trained, it functions as a general-purpose slider, enabling interpretable and fine-grained continuous control over various attributes. Moreover, by recombining the learned directions, the All-in-One Slider supports zero-shot manipulation of unseen attributes (e.g., races and celebrities) and the composition of multiple attributes. Extensive experiments demonstrate that our method enables accurate and scalable attribute manipulation, achieving notable improvements compared to previous methods. Furthermore, our method can be extended to integrate with the inversion framework to perform attribute manipulation on real images, broadening its applicability to various real-world scenarios. The code and trained model will be released at: https://github.com/ywxsuperstar/KSAE-FaceSteer.
Introduction
With the widespread emergence of diffusion models (Wu et al. 2023; Podell et al. 2023), text-to-image (T2I) generation has advanced significantly in recent years. Despite their impressive performance, a major challenge remains: providing users with continuous and fine-grained control functions over specific image attributes, which is essential for users to adjust generated images to obtain the desired results. As illustrated by the example in Figure 1(1), traditional prompt engineering approaches, such as simply appending phrases like “with a big smile” to the base prompt, “A photo of a gentle Thai man with layered hair, in a warm-toned studio”, often result in coarse-grained and rigid manipulations. Such textual variations not only prevent users from specifying the intensity of the smile but also inadvertently affect many unrelated attributes, such as the subject’s hairstyle, pose, and identity, leading to suboptimal outcomes.
Figure 1: Our All-in-One Slider shows advantages in: (1) Fine-grained and continuous control over desired attribute, without affecting other attributes (e.g., subjects’ identity and appearance). (2) Combination of multiple facial attributes (e.g., smile and age) for consistent and conflict-free transformations. (3) Zero-shot generalization to unseen attributes, without multiple and cumbersome training processes.
To enable attribute manipulation, early studies focused on exploring the representation properties of generative models and identifying the semantic directions corresponding to attributes. These efforts have spanned different representation spaces, including the UNet bottleneck layer (Kwon, Jeong, and Uh 2022; Park et al. 2023), text embedding space (Gandikota et al. 2024; Baumann et al. 2025; Hertz et al. 2022), noise space (Dalva and Yanardag 2024), weight space (Dravid et al. 2024; Gandikota et al. 2024), and representation in GAN-based architecture (Li, Hou, and Loy 2024; Tov et al. 2021). However, these methods often suffer from semantic entanglement (Kwon, Jeong, and Uh 2022; Park et al. 2023), lack clear attribute correspondence (Dalva and Yanardag 2024), or rely on the addition of phrases with limited continuity of attribute manipulation (Hertz et al. 2022). Some other studies (Li, Hou, and Loy 2024; Tov et al. 2021) achieve precise control, but their methods are designed for the GAN spaces, restricting the generalization to the powerful diffusion models. As a result, these methods fail to achieve both fine-grained and continuous attribute manipulation within a single framework.
Recent studies (Dravid et al. 2024; Gandikota et al. 2024; Baumann et al. 2025) have made efforts in achieving continuous attribute manipulation (e.g., gradually intensifying one’s smile or progressively increasing the degree of aging). Some methods  (Dravid et al. 2024; Gandikota et al. 2024) rely on domain-specific LoRA training to achieve latent space disentanglement. Others (Baumann et al. 2025) introduce attribute-wise training, which enables continuous attribute manipulation but requires attribute-wise supervision and heavily relies on labeled or paired data. As a result, these methods typically require separate training for each attribute and follow a One-for-One design paradigm. In their pipeline, each attribute is handled in isolation via a dedicated slider module or a specific optimized representation direction, as shown in Figure 2(1). When manipulating multiple attributes, this One-for-One paradigm requires multiple training runs, leading to significant time and computational costs and limiting their scalability and applicability in real-world scenarios.
Figure 2: (1) Existing One-for-One slider methods require training a specific slider module for each attribute. (2) Our All-in-One slider only needs training once to obtain a unified representation for different attributes, supporting the flexible manipulation of multiple diverse attributes.
Facing the limitations of the One-for-One paradigm, we propose an All-in-One Slider framework, as illustrated in Figure 2(2), which uses a single lightweight module to achieve flexible control over multiple attributes. In order to achieve generalized control of multiple attributes within a single module, a key challenge lies in disentangling these attributes. Once disentangled, the intensity of a target attribute can be easily controlled without affecting others, therefore enabling generalized and precise attribute manipulation. Inspired by the recent studies of sparse autoencoder (SAE) in LLMs (Karvonen et al. 2024; Huben et al. 2023), we adopt their idea of “break it down to build it up” and propose an Attribute Sparse Autoencoder, which is implemented on the text embeddings used in the image generation process. It first encodes each embedding into many fundamental components in a unified latent space for diverse attributes. By controlling the activation of these components, the desired attributes are then semantically created through decoding the activated ones. Specifically, we collect a set of commonly used textual prompts in portrait generation as our training dataset and encode them into the text embeddings. Subsequently, the Attribute Sparse Autoencoder is trained to perform auto-encoding on these embeddings and activate as sparsely as possible, in which the sparsity is the key to attribute disentanglement. During inference, we can control the activation of corresponding components in the latent space. By directly manipulating the coefficient of activated components, it naturally supports continuous and fine-grained control over desired attributes.
Through operating in this sparsely activated latent space of attributes, our All-in-One Slider can achieve decoupled control of multiple attributes through a single module, without the need for any specific training for each attribute. Moreover, as a unified attribute manipulation framework, our All-in-One Slider shows scalability and generalization to unseen attributes in T2I generation. Both qualitative and quantitative experiments are conducted to demonstrate the effectiveness of our All-in-One Slider for continuous and fine-grained manipulation of attributes.
The main contributions of this paper are as follows:
  • We propose a lightweight Attribute Sparse Autoencoder to achieve precise and continuous manipulation of individual or multiple attributes of image generation.
  • We propose the All-in-One Slider framework, enabling unified manipulation covering a wide range of diverse attributes, without requiring multiple training sessions.
  • Benefiting from constructing a unified latent space of diverse attributes, for the first time, our method shows the zero-shot continuous control for unseen attributes.
Related Work
Attribute Manipulation in T2I Diffusion Models
Controllability has become a central goal in T2I diffusion models, enabling users to steer generation toward desired contents or styles. Early approaches, such as ControlNet (Zhang, Rao, and Agrawala 2023; Mou et al. 2024), introduce spatial guidance by conditioning on structural cues like edges, poses, or depth maps. These methods are effective at preserving geometric alignment but are inherently limited when it comes to controlling high-level semantic attributes—such as age, expression, or makeup—which are not explicitly represented in these inputs.
To achieve fine-grained semantic control beyond spatial guidance, researchers explored semantic manipulation in latent spaces (e.g., noise or $\mathcal{H}$-space). Methods like SEGA (Brack et al. 2023) and Boundary Guidance (Zhu et al. 2023) enable editing without training, but lack fine-grained continuity or compositional control over semantic attributes. Later works shifted toward manipulating the text embedding space to achieve more direct semantic control. Prompt-based methods (Hertz et al. 2022; Brooks, Holynski, and Efros 2023) intervene through text prompts and attention maps, but often suffer from discrete or coarse edits. Meanwhile, editing methods like Imagic (Kawar et al. 2023) and ITI-Gen (Zhang et al. 2023) focus on aligning generation with user-provided images or preferences, but rely on image-text pairs and lack scalability for general attribute manipulation. Recent approaches, like ConceptSlider (Gandikota et al. 2024) and AttributeControl (Baumann et al. 2025), seek to enable continuous control by learning attribute vectors with paired data. However, they still require training a dedicated slider for each attribute, which limits scalability to unseen attributes.
To address these existing limitations, we propose a unified framework that discovers an interpretable and disentangled attribute latent space. This design enables smooth and scalable manipulation across multiple facial attributes without requiring retraining or paired labels, providing precise, continuous control in a plug-and-play manner.
Figure 3: An overview of our All-in-One slider’s framework. Stage 1: Unsupervised training of Attribute Sparse Autoencoder, which takes a token embedding obtained from the residual streamer in the text encoder as input and aims to reconstruct it with sparse features. Stage 2: Applying the trained Attribute Sparse Autoencoder to flexibly manipulate specific attributes during the image generation process.
Sparse Autoencoders (SAEs)
Sparse autoencoders (Ng et al. 2011; Olshausen and Field 1997; Bengio, Courville, and Vincent 2013) have been proven effective in learning interpretable representations in unsupervised settings (Coates, Ng, and Lee 2011). Among them, K-Sparse Autoencoders (K-SAE) (Makhzani and Frey 2013) enforce sparsity by allowing only the top-$k$ activations to activate, promoting the emergence of semantically meaningful units.
Some studies in the context of large language models (Gao et al. 2024; Cunningham et al. 2023) introduce SAEs in capturing discrete semantic concepts. Recent studies leverage the SAEs to interpret and analyze the internal mechanisms of diffusion models. SAeUron (Cywiński and Deja 2025) uses SAEs for concept unlearning by suppressing undesired features across diffusion steps. Diffusion Lens (Toker et al. 2024) decodes intermediate text encoder activations to reveal semantic representations. Unpacking SDXL Turbo (Surkov et al. 2025) decomposes transformer blocks to expose interpretable, causal factors across style, structure, and semantics. However, these works primarily focus on model interpretability and causal tracing. In this work, we extend K-SAE into a controllable face attribute manipulation framework that serves as a lightweight and scalable solution, enabling precise attribute control without retraining the base model.
Method
To break away from the traditional One-for-One paradigm, where each attribute requires a dedicated module or control path, we propose the All-in-One Slider framework. The core architecture of our framework is a highly parameter-efficient module, the Attribute Sparse Autoencoder. It achieves attribute disentanglement modeling and constructs a unified attribute latent space. This design enables precise, composable, and generalizable control over multiple visual attributes without retraining or attribute-specific supervision. As shown in Figure 3, the All-in-One Slider contains two key stages: (1) learning a latent representation from intermediate text embeddings, and (2) manipulating the diffusion generation process by modifying these latent codes. The full procedure is outlined in Algorithm 1.
Preliminary: Diffusion-based T2I Generation
With the advent of DDPM (Ho, Jain, and Abbeel 2020), diffusion models (Saharia et al. 2022; Nichol et al. 2021) have become the dominant technology in visual generation, especially for many state-of-the-art T2I generation systems. Recent T2I diffusion models typically follow the latent generation process (Rombach et al. 2022) and consist of three main components: a VAE (Kingma and Welling 2013), a text encoder, and a diffusion generator. The VAE compresses images into a latent space and decodes the generated latent back to pixel space. The text encoder, such as CLIP (Radford et al. 2021), encodes textual prompts into embeddings. The diffusion generator then takes a noisy latent as input and captures prompt information through cross-attention with the text embeddings, iteratively denoising the noisy latent towards a final output, enabling text-to-image generation.
Algorithm 1 Train Slider & Attribute Manipulation
Input: Text prompt $x_{\text{txt}}$, attribute text $A$, scalar $\lambda$
Output: Manipulated image II
1: Initialize ENC and DEC
[…]
3:
4:STAGE 1: Attribute Disentanglement Training
5:for $i\leftarrow 1$ to NN do
6:  $x\leftarrow\mathcal{D}$
7:  $z_{\text{ALS}}\leftarrow\text{Top-k}(\text{ReLU}(W_{\text{enc}}(x-b_{\text{pre}})+b_{\text{enc}}))$
8:  $\hat{x}\leftarrow W_{\text{dec}}z_{\text{ALS}}+b_{\text{pre}}$
9:  $\mathcal{L}=\|x-\hat{x}\|^{2}_{2}+\alpha\mathcal{L_{\text{aux}}}$ # $\mathcal{L}_{\text{aux}}$: auxiliary loss
10:end for
11:
12:STAGE 2: Attribute Manipulation
[…]
17:Output: II
Attribute Disentanglement Training
To achieve encoding and memorization of as many attributes as possible with limited parameters, we introduce the Attribute Sparse Autoencoder, which contains an encoder and a decoder. For training, we collect numerous text embeddings from the text encoder of a pretrained diffusion model. The encoder part of the Attribute Sparse Autoencoder maps these embeddings into a high-dimensional latent space (referred to as Attlatentspace in the following). Once trained, this Attlatentspace can serve as a unified attribute space, where attributes with different semantics sparsely activate a minimal and distinct set of neurons, thereby enabling disentangled modeling of diverse attributes. The decoder part then reconstructs the original embeddings from the latent representations. To encourage sparsity and disentanglement in the Attlatentspace, we draw inspiration from SAEs (Makhzani and Frey 2013; Gao et al. 2024; Karvonen et al. 2024) and apply a top-$k$ activation strategy to sparsely activate the attribute latent space. The overall training procedure consists of three major steps: embedding extraction, sparse encoding, and sparse reconstruction, as detailed below.
Embedding Extraction
The text encoder of a pretrained diffusion model (e.g., SDXL) consists of a stack of transformer blocks $\{B_{1},\ldots,B_{L}\}$. We extract the hidden state $x$ from an intermediate block $B_{l}$ within this stack. Specifically, we extract features from both SDXL text encoders, concatenate them to form the final joint representation $x$, and feed it into the Attribute Sparse Autoencoder.
Sparse Encoding
We implement sparsity using a top-$k$ selection mechanism. The encoder ENC maps the extracted embedding $x$ into the Attlatentspace $z_{\text{ALS}}$ via a linear transformation, ReLU activation, and top-$k$ selection:
$z_{\text{ALS}}=\text{Top-}k\left(\text{ReLU}(W_{\text{enc}}(x-b_{\text{pre}})+b_{\text{enc}})\right).$ (1)
Here, retaining only the top-$k$ neurons with the highest activation yields a compact and disentangled representation, where each attribute is associated with a distinct subset of neurons. This sparsity facilitates precise attribute manipulation during generation. The decoder then reconstructs $\hat{x}$ as:
$\hat{x}=W_{\text{dec}}z_{\text{ALS}}+b_{\text{pre}}.$ (2)
Sparse Reconstruction
We train Attribute Sparse Autoencoder to minimize the reconstruction loss:
$\mathcal{L}_{\text{mse}}=\|x-\hat{x}\|^{2}_{2}.$ (3)
However, a known limitation of sparse autoencoders is the prevalence of dead neurons—latent units that often remain inactive throughout training. These unused neurons reduce the expressiveness of the learned space and hinder future composability. To address this, we adopt an auxiliary mechanism inspired by (Gao et al. 2024). At each training step, we compute the residual $r=x-\hat{x}$ and identify a set of $k_{\text{aux}}$ least-activated neurons. The corresponding activations, denoted as $\hat{z}_{\text{ALS}}$, are used to reconstruct the residual:
$\hat{z}_{\text{ALS}}=\text{Top-}k_{\text{aux}}\left(\text{ReLU}(W_{\text{enc}}(x-b_{\text{pre}})+b_{\text{enc}})\right).$ (4)
We compute the auxiliary reconstruction $\hat{r}$ as $\hat{r}=W_{\text{dec}}\hat{z}_{\text{ALS}}$, and define the auxiliary loss as: $\mathcal{L}_{\text{aux}}=\|r-\hat{r}\|^{2}_{2}.$ The final training objective thus becomes:
$\mathcal{L}=\mathcal{L}_{\text{mse}}+\alpha\mathcal{L}_{\text{aux}}.$ (5)
This loss encourages even initially inactive neurons to explain parts of the residual, ensuring broader semantic coverage in the latent space. After training, each dimension in the sparse latent space corresponds to a specific, disentangled facial attribute. This latent space serves as the basis for attribute manipulation, as detailed in the next section.
Attribute Manipulation
Once the Attribute Sparse Autoencoder is trained, it can be employed as an “Attribute Manipulation Slider”, enabling control over specific attributes within the unified Attribute Latent Space of the text encoder. Given an attribute $A$ that we expect to manipulate, we extract its corresponding text embedding $x_{A}$ and feed it to the trained Attribute Sparse Autoencoder to identify the latent direction associated with the attribute. We then modify the input prompt embedding $x$ by adding this direction scaled by a scalar $\lambda$, as follows:
$x_{\text{manipulated}}=x+W_{\text{dec}}(\lambda\times\text{ENC}(x_{A})),$ (6)
where $\lambda$ controls the strength of the steering. This modification of the embedding $x_{\text{manipulated}}$ is then used to condition the generation process, enabling precise manipulation of specific attributes without disrupting the overall quality or structure of the generated image. Importantly, this manipulation operates entirely within the learned sparse latent space, supporting zero-shot, composable editing across a wide range of diverse attributes. This unified mechanism underpins the scalable and flexible attribute control explored.
Experiments
Experimental Setup
Implementation Details
We employ SDXL (Podell et al. 2023) to generate images using the DPM-solver sampler over 50 sampling steps with classifier-free guidance (Ho and Salimans 2022) of 7.5. All the compared methods are implemented following their default configurations available from their official repository. Besides, we edit real images using SDXL-Turbo (Sauer et al. 2024) with the Renoise inversion method (Garibi et al. 2024). We extract features from the 10th and 29th layers of the two text encoders in SDXL, which contain 12 and 32 transformer layers, respectively. We train our Attribute Sparse Autoencoder with 52,000 text samples (52 controllable facial attributes, and 1000 samples per attribute). The details of construction and Attribute Sparse Autoencoder parameters used for the experiments are included in the supplementary material.
Quantitative Metrics
Traditional metrics, such as the CLIP score (Hessel et al. 2021), are limited in their ability to evaluate fine-grained or localized attribute manipulations. These methods typically measure overall semantic similarity, which may obscure whether specific target attributes were correctly edited. To more comprehensively assess image editing quality and manipulation capabilities, we adopt two complementary metrics specifically designed to capture the nuanced effects of facial attribute manipulation. We use the multimodal large language model, Qwen2.5-VL-Instruct (Wang et al. 2024), to evaluate the semantic alignment between the intended textual prompt (e.g., “smile”, “old”, and “makeup”) and the generated image, termed as Qwen Score (QS) (Han et al. 2025). In contrast to traditional metrics, Qwen demonstrates superior reasoning ability and aligns more closely with human judgment. The QS evaluates the edited image relative to the original image based on three criteria: Action Fidelity (the alignment with the edit prompt), Identity Preservation (the consistency of the subject’s identity), and Visual & Anatomical Coherence (the naturalness of the modifications). The average of these three scores is used as the final metric. ID Consistency Score (IS): We further measure the identity preservation between the source image and the edited result using an ID consistency score based on a face recognition model (i.e., Arcface (Deng et al. 2019)). This quantifies how well our method preserves the core identity features despite attribute manipulations. Identity preservation is especially important in real-image editing scenarios, where users expect natural-looking changes without sacrificing personal identity. Higher QS and IS both indicate better editing performance, reflecting more accurate attribute manipulation and stronger identity preservation.
Text-to-Image Attribute Manipulation
To evaluate the ability of All-in-One Slider to manipulate attributes, we present qualitative results in Figure 4. The wide range of visual controls, which covers both semantic attributes (e.g., smile, age, makeup) and physical accessories (e.g., eyeglasses, hat), demonstrates the expressive power and controllability of our method. Each manipulated image preserves the identity and background of the original input, while exhibiting precise and consistent attribute changes. This controllability is enabled by the sparsity and disentanglement of directions in the latent space of Attribute Sparse Autoencoder, where each attribute corresponds to an independent and well-isolated latent direction. This allows precise manipulation without affecting unrelated visual features. In contrast, traditional methods (Baumann et al. 2025; Gandikota et al. 2024) rely on individual sliders per attribute, which inherently restricts the controllable attribute space as each new attribute necessitates its own slider. Our unified framework overcomes this constraint, enabling efficient manipulation across a broader range of attributes.
Figure 4: Qualitative results of face attribute manipulation. Our All-in-One slider can perform both fine-grained semantic edits (e.g., smile, makeup, and age) and physical changes (e.g., eyeglasses, hat, hair style, and skin tone).
Figure 5: Qualitative results of compositional multi-attributes manipulation. Our All-in-One slider achieves coherent manipulation while preserving the original identity.
Multi-Attribute Manipulation
To evaluate the compositional controllability of our method, we perform multi-attribute manipulation experiments as shown in Figure 5. Starting from a base image, we simultaneously apply multiple attributes such as Smile, Makeup, Hat, and Old. The edited images show a visually coherent and semantically consistent blend of attributes without artifacts or conflicts, demonstrating the model’s ability to naturally combine multiple attributes while still preserving the original identity. Similar compositional behavior is also observed in Figure 1 (2), where various attribute combinations (e.g., Smile, Old, Eyeglasses) are applied, yielding realistic and conflict-free results across identities. This compositional behavior stems from our unified latent space of Attribute Sparse Autoencoder, where different attribute directions are effectively disentangled yet compatible. Since all manipulations are represented as sparse vectors within a shared embedding space, their combinations remain semantically coherent and do not interfere with each other.
Quantitative Analysis
We quantitatively evaluate the effectiveness of our method in both single-attribute (Sng) and multi-attribute (Mul) manipulating tasks, using QS and IS as evaluation metrics. We compare against two One-for-One baselines: ConceptSlider (Gandikota et al. 2024) and AttributeControl (Baumann et al. 2025). Results are summarized in Table 1. In the single-attribute setting, our method achieves competitive performance, obtaining the highest IS for the Old attribute and second-best scores for Smile and Makeup in both QS and IS. In contrast, in the multi-attribute setting—where compositional complexity and entanglement pose greater challenges—our method outperforms the baselines. It achieves the highest QS and IS scores across all combinations (Old+Smile, Old+Makeup, and Smile+Makeup). Notably, for the challenging Old+Makeup case, our approach achieves QS of 4.4281 and IS of 0.6277, surpassing both baselines by a clear margin.
These results highlight the core strength of our approach: the ability to perform disentangled, compositional steering while preserving identity. Unlike prior methods that often sacrifice attribute fidelity for identity preservation, our method strikes a better balance, maintaining both attribute accuracy and subject identity in both single- and multi-attribute manipulation tasks. The superior performance, particularly in multi-attribute cases, stems from our disentangled Attribute Latent Space, which enables conflict-free, compositional control over multiple attributes without compromising the subject’s core identity.
Set Method Old Smile Makeup
QS IS QS IS QS IS
Sng CSlider 3.7941 0.4336 4.1437 0.4962 4.5424 0.6529
AttCtrl 4.0392 0.6005 4.3954 0.6952 4.2680 0.6044
Ours 4.0490 0.7155 4.2647 0.6366 4.2908 0.7423
Old+Smile Old+Makeup Smile+Makeup
Mul CSlider 4.1503 0.4993 3.8006 0.5224 4.0588 0.4786
AttCtrl 3.6667 0.3755 4.0555 0.6347 4.2484 0.5145
Ours 4.2124 0.6882 4.4281 0.6277 4.2973 0.6351
Table 1: Quantitative comparisons across single and multiple attributes editing. “QS” denotes the semantic alignment scored by Qwen, and “IS” denotes the identity consistency score. The best and second-best results are highlighted with bold and underline, respectively.
Model Generalization
We further investigate the architectural generalization of our method by applying it to two different diffusion models: Stable Diffusion v1.4 (SD v1.4) and SDXL-Turbo. In Figure 6 (1), progressively increasing the control strength $\lambda$ yields fine-grained variations in makeup intensity, demonstrating effective attribute modulation within the SD v1.4. Figure 6 (2) further validates our method on SDXL-Turbo, where semantic manipulation of Smile attribute is applied with high fidelity and strong identity preservation. These results highlight the robustness and transferability of our approach across distinct models.
Reference-Based Image Editing
To enable high-fidelity editing on real images, we integrate our method with the ReNoise inversion framework (Garibi et al. 2024), which reconstructs latent representations from real inputs. As illustrated in Figure 7, we conduct editing on two representative attributes: Smile (top) and Old (bottom). In the first row, when adding the Smile, our method clearly avoids unnatural facial deformation (e.g., overly rounded cheeks for AttControl). In the second case, it successfully preserves accessories like eyeglasses during the addition of complex attributes such as Christmas elements with the addition of old. Compared to AttributeControl, our approach better maintains critical identity features, which is primarily attributed to our Attribute Sparse Autoencoder-based decoupling strategy, which minimizes unintended attribute entanglement and improves the overall stability and locality of edits via sparse semantic modulation.
Figure 6: Qualitative results on SD v1.4 (1) and SDXL-Turbo (2) showing the effectiveness of our method in attribute modulation (e.g., makeup intensity and smile generation) across U-Net-based architectures.
Figure 7: Real image editing. We apply our method to edit Smile (top) and Old (bottom) attributes using the ReNoise inversion framework (Garibi et al. 2024). Compared to AttributeControl, our method better preserves identity details like eyeglasses and facial structure.
Figure 8: Continuous zero-shot generalization to different racial attributes. The model progressively manipulates facial features across varying strengths of the racial attribute ($\lambda=0.3,0.4,0.5$).
Zero-shot Generalization Ability
Although our All-in-One Slider is trained on a fixed set of facial attributes, it demonstrates strong zero-shot generalization to new racial and identity attributes not encountered during training. As shown in Figure 8, the model can continuously modify facial attributes, including races (e.g., African, Korean, Chinese, Hispanic, and Indian) and celebrity-like identities (e.g., Barack Obama and Einstein in Figure 1 (3)), without the need for additional fine-tuning or supervision.
This generalization ability arises from our attribute decomposition strategy, where each facial attribute is learned as a set of fine-grained, disentangled components within the shared Attlatentspace. The pretrained text encoder has built the alignment of generalized objects and their attribute correspondences. For objects (e.g., races and celebrity names) that have not been seen during training but are memorized by the generative model itself, their text embedding can still induce adaptive combinatorial activation of attribute components within Attlatentspace to reconstruct corresponding complex combinations of key attributes. This ability to perform zero-shot and open-vocabulary manipulation of attributes further underscores the versatility of our approach for controllable image synthesis.
Ablation Study
We conduct ablation studies on two key components of our method: the manipulated text encoder layers and the steering strength used during attribute control.
Layer Selection
We evaluate manipulating performance across different combinations of transformer layers, spanning from intermediate (8/28) to deep (10/28) configurations in SDXL’s dual text encoders. These layers have the most significant impact on generated images, covering both shallow and deep semantic regions. As shown in Table 2, deeper layers generally yield higher Qwen scores (semantic score), while identity scores tend to decline. The 10/28 combination achieves the best overall balance, indicating that mid-level layers are more suitable for simultaneously preserving semantic intent and identity. This clearly reflects the practical importance of choosing appropriate feature granularity for localized and identity-sensitive editing in practice.
Effect of Manipulating Strength $\lambda$
Figure 9 shows how varying $\lambda$ affects the trade-off between manipulation strength and identity retention. Lower values (e.g., $\lambda=0.15$) lead to under-editing, whereas higher values (e.g., $\lambda=0.30$) result in stronger attribute expression. We observe that $\lambda=0.20$ achieves a favorable balance, yielding strong attribute alignment (QS) while maintaining high IS across all attributes. In particular, degradation is most severe on “old”, suggesting that aging transformations are more entangled with identity features. These results demonstrate the inherent antagonistic effect between attribute change and identity preservation in the attribute manipulation.
X/Y Old Smile Makeup Avg
QS IS QS IS QS IS QS IS
8/28 3.961 0.672 4.216 0.537 4.196 0.696 4.124 0.635
9/30 4.062 0.631 4.193 0.617 4.177 0.759 4.144 0.669
10/24 4.023 0.722 4.245 0.668 4.288 0.765 4.185 0.718
10/28 4.049 0.716 4.265 0.637 4.291 0.742 4.202 0.698
Table 2: Evaluation for attributes old, smile, and makeup across different Manipulation Layer (X/Y) of SDXL’s dual text encoders, where X is the selected layer index in the first encoder and Y in the second.
Figure 9: Effect of manipulating strength ($\lambda$) on attribute manipulation (Qwen score) and identity preservation (ID consistency) for different attributes (old, smile, makeup).
Conclusion
We propose the All-in-One Slider, a sparse and lightweight framework for controllable face image generation using the Attribute Sparse Autoencoder. Unlike previous approaches that rely on attribute-specific training, our approach introduces a unified and disentangled latent space, enabling fine-grained, continuous manipulation of multiple facial attributes while preserving identity and visual fidelity. Extensive experiments demonstrate strong performance of our method in both single- and compositional-attribute manipulation, as well as zero-shot generalization to unseen attributes. All-in-One Slider presents a new and unified paradigm for continuous attribute control of image generation, offering scalable, interpretable, and flexible advantages. It is promising to inspire more work to achieve continuous attribute control with low training and parameter cost based on the unified space.
Supplementary Material
1    Implementation Details
Training Details
We use SDXL as the base generative model and extract intermediate text embeddings from its 11th (of 12) and 29th (of 32) transformer layers blocks (layer.10 and layer.28) across the dual encoders to train the Attribute Sparse Autoencoder. The activations are computed over prompts covering 52 attributes, each augmented with diverse identity-aware prefixes to encourage broad semantic coverage. The dimension of text embedding is 2048 (concatenation of 768 and 1280). The autoencoder has a latent dimension of 32,768 with an expansion factor of 16. It is trained using the Adam optimizer with a learning rate of $4\times 10^{-4}$ and a batch size of 4096 latent tokens per step. The total training budget is 400 million tokens, corresponding to approximately 97,656 training steps. During attribute selection, we retain the top-$k=128$ most activated dimensions. To mitigate overfitting and encourage broader attribute coverage, we introduce an auxiliary selection mechanism with $k_{\text{aux}}=256$ and apply a regularization coefficient of $\alpha=0.1$ to increase its influence during training. All experiments were performed using a single NVIDIA GeForce RTX 4090 GPU with 24 GB of VRAM.
Training Prompt Construction
To expose the Attribute Sparse Autoencoder to a wide variety of visual concepts, we construct a diverse set of text prompts associated with over 52 semantic attributes (e.g., age, arched eyebrows, wearing necklace). For each attribute, we manually define several textual expressions that describe different attribute states (e.g., “old man”, “very old person”, “young man”, “very young person”).
To enrich visual diversity and reduce identity bias, we prepend each prompt with descriptive prefixes drawn from a fixed pool. These include phrases like “a photo of a”, “a photo of a beautiful”, as well as identity-related descriptors such as “a photo of a black/white/asian/arab/caucasian/african-american”, and their compositional forms (e.g., “a photo of a beautiful asian”).
We construct the full prompt set by exhaustively combining all prefixes and prompt templates for each attribute. If the total number of combinations is less than 1000, we retain all; otherwise, we randomly sample 1000 combinations using different random seeds to ensure diversity across prompt sets. This results in a balanced and diverse prompt distribution per attribute.
Attribute List
We consider a total of 52 facial attributes and use their corresponding text embeddings to train the Attribute Sparse Autoencoder. These attributes span appearance, expression, hairstyle, and accessory-related features. The full list is as follows: age, angry, arched-eyebrows, bags-under-eyes, bald, bangs, big-lips, big-nose, birthmarks, black-hair, blond-hair, brown-hair, bushy-eyebrows, chubby, colorful-outfit, double-chin, elegant, eye-size, eyeglasses, fitness, freckled, frowning, gray-hair, groomed, high-cheekbones, makeup, moles, mouth-slightly-open, multicolored-hair, mustache, narrow-eyes, oval-face, pale-skin, pierced, ponytail, pouting, receding-hairline, rosy-cheeks, sad, scarred, sideburns, smile, surprised, tattooed, tiredness, wavy-hair, wearing-earrings, wearing-hat, wearing-lipstick, wearing-necklace, wearing-necktie, width
Image Editing Evaluation Criteria
In the evaluation of image edits, we introduce Qwen Score (QS) to assess changes in motion or facial expression. This is done using a 5-point scale across three dimensions: Action / Expression Fidelity, Identity Preservation, and Visual & Anatomical Coherence. This system is based on the scoring framework provided by the Qwen2.5-VL-Instruct model. The criteria are as follows:
Expression Fidelity captures how faithfully the edited image reflects the target expression described in the instruction. A high score reflects a precise match in terms of gesture, intensity, and orientation. In contrast, lower scores are triggered by absent or incorrectly rendered expressions, partial changes (e.g., only some facial muscles involved), or misaligned pose dynamics.
Identity Preservation evaluates whether the subject remains recognisable and consistent across the original and edited images. Perfect preservation includes consistent facial structure, hairstyle, and clothing. Score degradation occurs when noticeable drifts—such as altered facial features, skin tone, or missing accessories—make the subject appear different or less identifiable.
Visual & Anatomical Coherence focuses on the realism and structural consistency of the rendered image. High scores reflect anatomically correct bodies, natural lighting and shadows, and coherent textures. Lower scores are given in the presence of visual artifacts, anatomical distortions, or rendering inconsistencies such as cut-out edges, lighting mismatch, or unrealistic joint positions.
Each image pair (before and after editing) is evaluated in terms of three dimensions: Action Fidelity, Identity Preservation, and Visual & Anatomical Coherence, each scored on a 5-point scale. In addition to the numerical scores, a brief reasoning (within 20 words) is provided to justify the evaluation results. As an example, we present the evaluation of an edited image, for which the Qwen model assigned scores of 4 for Action Fidelity, 5 for Identity Preservation, and 4 for Visual & Anatomical Coherence (see Figure 10). The edited image shows a successful transition from a neutral expression to a clear, natural-looking smile, while preserving the head pose and facial features. Minor inaccuracies in the smile’s angle or intensity lead to an Action Fidelity score of 4. The subject’s identity is perfectly preserved, justifying a score of 5 for Identity Preservation. The image also maintains visual and anatomical coherence, with only slight imperfections, resulting in a score of 4 for Visual & Anatomical Coherence. According to Qwen’s analysis, the brief explanation for this result is: “The subject’s expression has changed to a smile, but the overall pose remains the same”.
Figure 10: Attribute manipulation results for Smile. The image shows the subject before and after the expression modification.
2    Additional Experimental Results
Attribute Manipulation
We present additional qualitative results demonstrating our method’s ability to manipulate a wide range of facial attributes in an identity-preserving manner, as illustrated in Figures 11 to 20. Each figure below shows a series of faces in rows, where the leftmost column is the original input image, and the rightward images represent progressive edits along a specific semantic attribute.
3    Prompt Information
For reproducibility, we provide the text prompts and associated random seeds used to generate the images in the appendix figures.
Prompts for Main Figures
Below are the prompts used to generate the images shown in the main figures of the paper.
Figure 1: (1) A photo of a gentle Thai man with layered hair, in a warm-toned studio. 33094
Figure 1: (2) A photo of a honest Dutch man in a white button-up shirt, with a blurred background. 27454
Figure 1: (3) A high-quality portrait of a man in a tuxedo, with clear, detailed facial features. 33673
Figure 4: Prompt: A close-up of a woman wearing a red scarf, looking thoughtful. 31796
Figure 5: A photo of a honest Latin American talented woman. 33819
Figure 6: (1) A woman with medium-length blonde hair, pale skin, and soft blue eyes.33819
Figure 6: (2) A photo of a peaceful Middle Eastern woman. 22436
Figure 7: (1) A man wearing glasses while holding a banana.
Figure 7: (2) A man in front of a Christmas tree with his dog.
Figure 8: A photo of a woman wearing a red scarf. 42
Prompts for Appendix Figures
Below are the prompts used to generate the images shown in the appendix.
Figure 11 prompts:
(1) A photo of a charming Singaporean man with clean-cut hair, with studio backdrop. 16727
(2) A photo of a gentle Thai man with layered hair, in a warm-toned studio. 26662
(3) A photo of a composed Swedish man with short blond hair, with a blurred background. 38836
(4) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 20664
(5) A photo of a peaceful Middle Eastern woman. 42
(6) A photo of a composed Moroccan man wearing a woolen scarf, with a blurred background. 29425
(7) A photo of a kind Turkish adult woman. 26281
(8) A photo of a honest Dutch man in a white button-up shirt, with a blurred background. 32119
Figure 12 prompts:
(1) A man in a white button-up shirt, with soft natural lighting on his face. 3203
(2) A photo of a refined Italian man wearing a checkered blazer, with soft window light. 22088
(3) A photo of a calm South African man wearing round glasses, with a soft gray backdrop. 6839
(4) A photo of a radiant American man wearing a casual sweater, with elegant lighting. 28031
(5) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 46112
(6) A photo of a serene Danish man with round glasses, near a sunlit window. 37614
(7) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows. 7242
(8) A photo of a thoughtful Latin American woman in a white button-up shirt, with a blurred background. 26731
Figure 13 prompts:
(1) A photo of a stylish Serbian person with long hair, in a candle-lit room. 42
(2) A photo of a refined Italian man wearing a checkered blazer, with soft window light. 48991
(3) A photo of a graceful British man with neatly combed hair, in a bright workspace. 33898
(4) A photo of a composed Swedish man with short blond hair, with a blurred background. 38623
(5) A photo of a stylish Belgian person with a fringe haircut, in a softly lit room. 44007
(6) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows. 10133
(7) A photo of a poised Greek person in a linen shirt, in a studio with soft shadows. 7242
(8) A photo of a composed Swedish man with short blond hair, with a blurred background. 38836
Figure 14 prompts:
(1) A photo of a honest Latin American talented woman. 33819
(2) A photo of a contemplative Pakistani woman with a patterned scarf, with natural daylight. 46948
(3) A photo of a charming Mongolian person wearing a fur-lined coat, in snowy weather. 2337
(4) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room. 6603
(5) A photo of a contemplative Pakistani woman with a patterned scarf, with natural daylight. 26926
(6) A photo of a unique Filipino woman. 48084
(7) A woman with short hair, looking at the camera. 13476
(8) A photo of a charming Taiwanese person with medium-length hair, in a sunlit studio. 26595
Figure 15 prompts:
(1) A portrait of a woman standing by the beach at sunset, with soft and natural lighting. 42
(2) A photo of a woman smiling in a cafe setting. 42
(3) A photo of a elegant and kind woman. 344980
(4) A photo of a beautiful woman standing by the beach at sunset. 42
(5) A woman with short hair, looking at the camera. 39743
(6) A photo of a radiant American man wearing a casual sweater, with elegant lighting. 28031
(7) A photo of a honest Polish woman in a white button-up shirt, with a soft gray backdrop. 29600
(8) A photo of a thoughtful Latin American woman in a white button-up shirt, with a blurred background. 26731
Figure 16 prompts:
(1) A photo of a serene Danish man with round glasses, near a sunlit window. 37614
(2) A photo of a serene Danish man with round glasses, near a sunlit window. 47741
(3) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 26529
(4) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 46112
(5) A photo of a stylish Serbian person with long hair, in a candle-lit room. 42
(6) A photo of a composed Swedish man with short blond hair, with a blurred background. 29070
(7) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 19257
(8) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 26125
Figure 17 prompts:
(1) A portrait of a woman standing by the beach at sunset, with soft and natural lighting. 42
(2) A woman with short hair, looking at the camera. 13476
(3) A woman with short hair, looking at the camera. 41815
(4) A woman with short hair, looking at the camera. 39743
(5) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room. 21553
(6) A photo of a sophisticated Russian woman with a pearl necklace, in a minimalist room. 6603
Figure 18: prompts:
(1) A photo of a honest Dutch man in a white button-up shirt, with elegant lighting. 43448
(2) A photo of a elegant Chinese man wearing a red scarf, with a blurred background. 26099
(3) A photo of a confident Turkish man wearing a floral crown, with natural daylight. 2606
(4) A photo of a cheerful Filipino man wearing a red scarf, with a soft gray backdrop. 21968
(5) A woman with short hair, looking at the camera. 13476
(6) A photo of a vibrant Indian man. 41026
(7) A photo of a serene American man wearing a floral crown, with a soft gray backdrop. 41682
(8) A photo of a honest Middle Eastern adult woman wearing a red scarf, in a studio. 49637
Figure 19 prompts:
(1) A photo of a beautiful woman. 45678
(2) A photo of a beautiful woman. 355667
(3) A photo of a wise European man wearing a casual sweater, in a sunny environment. 26509
(4) A photo of a elegant woman. 569988
Figure 20 prompts:
(1) A portrait of a woman standing by the beach at sunset, with soft and natural lighting. 42
(2) A close-up of a woman with a floral crown on her head. 43533
(3) A photo of a honest Latin American talented woman. 33819
(4) A close-up portrait of a indian woman. 1646
(5) A photo of a radiant Nepalese woman wearing a gold necklace, in natural light. 26268
(6) A photo of a gentle Thai man with layered hair, in a warm-toned studio. 31527
(7) A photo of a gentle Vietnamese man with short cropped hair, with a soft gray backdrop. 10368
(8) A photo of a serene Brazilian confident woman. 45433
Note that the prompts correspond to the images in order, starting from the top-left image and proceeding row by row (left to right, then top to bottom).
Figure 11: Continuous manipulation of the Smile attribute. Each row shows a subject transitioning from neutral to smiling, with identity and lighting well preserved.
Figure 12: Continuous manipulation of the Old attribute. Each subject’s image is progressively transformed to appear older, showing realistic aging effects such as wrinkles and hair graying, while maintaining overall identity.
Figure 13: Continuous manipulation of the Wavy Hair attribute. Each row shows a smooth transition with identity preserved.
Figure 14: Continuous manipulation of the Wearing Lipstick attribute. Each row shows a smooth transition with identity preserved.
Figure 15: Continuous manipulation of the Afro hairstyle attribute. Each row shows increasing Afro intensity from left to right, with identity and background preserved.
Figure 16: Continuous manipulation of the Bald attribute. Each row shows increasing Bald intensity from left to right, with identity and background preserved.
Figure 17: Manipulation for the Blond Hair attribute. Our method alters hair color while maintaining identity and realism.
Figure 18: Manipulation of the Eyeglasses attribute. Each pair shows the original image and its edited version with eyeglasses added.
Figure 19: Attribute manipulation results for Hat. Each pair shows the subject before and after adding a hat.
Figure 20: Attribute manipulation results across various appearance traits. Each pair shows the subject before (left) and after (right) editing, demonstrating the effect of adding attributes such as black hair, scar, necklace, beard, receding hairline, and multicolored hair. The results illustrate the model’s ability to generate realistic and identity-consistent modifications across a diverse set of visual changes.